PROBING CIRCUMSTELLAR MOLECULAR GAS IN THE STAR-FORMING REGION BD +40^{o} 4124

\underline{SHIYA WANG} and LESLIE W. LOONEY, Astronomy Department, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

As a well known cluster system that has simultaneous low and high-mass star formation, the BD +40^{o} 4124 region can provide an interesting laboratory for star formation studies. The cluster associated with BD +40^{o} 4124, a Herbig Be star at a distance of 1 kpc, includes two other optically visible young stars, V 1686 Cygni and V1318 Cygni, and several embedded IR sources. We present the BIMA interferometric images of CS J = 2 − 1, ^{13}CO J = 1 − 0, and C^{18}O J = 1 − 0 to probe the molecular gas distribution and morphology and derive the gas column density and molecular mass. By comparsion with near infrared observations, we identify more embedded sources showing ongoing star-forming activities and study the formation mechanism and interaction between the high mass star and low mass companions. Finally, the impact to general star formation will be discussed.